Using Cramer's rule, find Y and C in the following systems of equations
Y=C+I+G
C=25+6Y^(1/2)
I=16
G=14
1
Expert's answer
2020-11-27T13:11:49-0500

Y=C+I+G, so:

"Y = 25 + 6Y^{0.5} + 16 + 14,"

"(Y^{0.5})^2 - 6Y^{0.5} - 55 = 0,"

Y1^(0.5) = 11 or Y2^(0.5) = -5,

Y1 = 121, Y2 = 25,

C1 = 25 + 6×11 = 61,

C2 = 25 + 6×(-5) = -5, so this root will not apply.